OpenStudy (anonymous):

Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x)

OpenStudy (anonymous):

2.5x-10.5=64\[2.5x-10.5=64(0.5^{x}\]

OpenStudy (unklerhaukus):

2.5x–10.5=64(0.5^x)

OpenStudy (unklerhaukus):

So what does the table look like?

OpenStudy (unklerhaukus):

ok so you have f(x) = 2.5x-10.5 to the first table entry is f(2) = 2.5(2)-10.5 =

OpenStudy (unklerhaukus):

so simplify this and put the answer in the top right box, next find f(3) = 2.5(3)-10.5 = then f(4), f(5), and f(6)

OpenStudy (anonymous):

im still confused

OpenStudy (unklerhaukus):

f(x) = 2.5x-10.5 is a function of the variable x the column on the far left is values for x the first blank table entry is f(x) = 2.5x-10.5 where x=2 i.e. f(2) = 2.5(2) - 10.5 = 2.5 × 2 - 10.5 =

OpenStudy (unklerhaukus):

so continue simplifying this until you get some number, this number is f(2), i.e. the function f(x) evaluated at the point x=2

OpenStudy (unklerhaukus):

can you simplify \[2.5 × 2 - 10.5\] @kiing_sosa ?

OpenStudy (anonymous):

can i just get an awnser

OpenStudy (unklerhaukus):

start with the multiplication 2.5 x 2 = . . .

OpenStudy (anonymous):

5

OpenStudy (unklerhaukus):

f(2) = 2.5(2) - 10.5 = 2.5 × 2 - 10.5 = 5 - 10.5 = . . .

OpenStudy (unklerhaukus):

now the take-away

OpenStudy (unklerhaukus):

5 - 10.5 = . . .

OpenStudy (anonymous):

aye look man this aint wokin aoll i need is anwsers ill medel you and fan

OpenStudy (unklerhaukus):

i'm not here to provide answers,

OpenStudy (anonymous):

well help a little more

OpenStudy (anonymous):

Dude he did all the work its just 5-10.5!

OpenStudy (anonymous):

if you need to just get a calculator but try a little!